area of a sail. the area in square meters for a triangular sail is given by
A(x)=x^2+5x+6
A) find A(5).
b) If the height of the sail is x+3 meters, then what is the length of the base of the sail.

Answer by edjones(8007)

A(x)=x^2+5x+6
A) A(5)=5^2+ 5*5 +6
=25+25+6
=56m^2
.
B)
A=.5ab
56=.5(5+3)b I am assuming that x=5. If it doesn't then ignore this answer.
4b=56
b=14m
.
Ed